OpenStudy (anonymous):

HELP: Show that F(x)=x+x^-1 is one to one on the interval {1, infinity) and find the inverse of F^-1(x)!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

If you graph the function, you can see that the function approaches a line on the interval (1, infinity)

OpenStudy (anonymous):

yeah but how do you find the inverse of the formula??

OpenStudy (anonymous):

Switch x and y in the function and then solve for y. So y=x+x^-1 becomes x=y+y^-1 and then just solve the equation for y.

OpenStudy (anonymous):

probably would help to write it as \[f(x)=x+\frac{1}{x}=\frac{x^2+1}{x}\]

OpenStudy (anonymous):

then solve \[x=\frac{y^2+1}{y}\] for \(y\) it will require solving a quadratic equation, so could be a bit tricky

OpenStudy (anonymous):

you know how to do that, or you need help?

OpenStudy (anonymous):

yes!!! thank you!!!1 makes sense